OpenStudy (anonymous):

Did anyone do PRE-CALCULUS PART II, SECTION 2?

OpenStudy (anonymous):

Do you have a question you need help with?

OpenStudy (anonymous):

Let B = [-1 3 6 -3]. Find -4B.

OpenStudy (anonymous):

A. [-4 12 24 -12] B. [-3 1 4 -5] C. [4 -12 -24 12] D. [4 3 6 -3]

OpenStudy (anonymous):

@LegendarySadist

OpenStudy (anonymous):

\[\large \sf [(-4 \times -1) (-4 \times 3) (-4 \times 6) (-4 \times -3)] \]

OpenStudy (anonymous):

so answer c @LegendarySadist

OpenStudy (anonymous):

Correct